XYZ Corporation has a dividend payout rate of 25%. The firm's current earnings of $18.50 per share are expected to grow at an annual rate of 5%, and its cost of equity capital is 10%. What is the firm's expected "price" per share?

a.

$19.425

b.

$25.90

c.

$46.25

d.

$77.70
Answer #1

Dividend0 = 18.50 * 0.25 = 4.625

Div1 = 4.625 * (1 + 0.05) = 4.85625

Constant growth Dividend discount model

Price = Div1/(r - g)

Price = 4.85625/(0.1 - 0.05)

Price = $97.125

Can you please upvote? Thank You :-)